OpenStudy (anonymous):

let f(x)=(2/9)x^(3/2) and suppose that the line y=(5/2)x is parallel to the tangent of f(x) at x=a find the value of a. please show all steps!

OpenStudy (anonymous):

dydx return the value of gradient of the point . dy/dx f(x)=1/3x^(1/2) the second equation show gradient = 5/2 then, 1/3x^(1/2)=5/2 x=15²/2²

OpenStudy (kira_yamato):

So this means to say, at x = a, f'(x) = 5/2. Differentiating f(x) with respect to x, f'(x) = sqrt(x)/3 sqrt(x) = 15/2 x = 56.25